# Strada le Grazie 15, Verona, Italy

### v1

@@ Abstract
This study introduces the Iterative Refinement Process (IRP), a robust anomaly detection methodology tailored for high-stakes industrial quality control. The IRP systematically enhances defect detection accuracy by employing a cyclic data refinement strategy that iteratively removes misleading data points, thereby improving the model’s performance and robustness. We validate the effectiveness of the IRP using two benchmark datasets, Kolektor SDD2 (KSDD2) and MVTec-AD, covering a wide range of industrial products and defect types. Our experimental results demonstrate that the IRP consistently outperforms traditional anomaly detection models, particularly in environments with high noise levels. This study highlights the IRP’s potential to significantly enhance anomaly detection processes in industrial settings, effectively managing the challenges of sparse and noisy data.
Keywords:
Robust Anomaly Detection Self Supervised Learning Iterative Refinement Process.
@@ 1 Introduction
Anomaly Detection (AD) plays an indispensable role in quality control across a wide range of manufacturing industries, ensuring the integrity of materials such as marble [1], steel [2], and leather [3]. Accurately detecting items that are non-compliant with product specifications is crucial for maintaining product standards and consumer satisfaction. However, this task is fraught with challenges, primarily due to the diverse and complex textures of the materials, the rarity of defects, and the significant scarcity of accurately labeled data necessary for effective supervised learning. Anomalies typically occupy only a tiny portion of an image, making their detection difficult for human inspectors and automated systems. This challenge is compounded by the labor-intensive and error-prone nature of manually labeled training data, a process that becomes increasingly unsustainable in high-throughput manufacturing settings.
The evolution of anomaly detection in recent years has increasingly leaned towards unsupervised learning paradigms, prized for their ability to operate without extensively labeled datasets. These approaches, while innovative, presuppose the existence of pristine training data, free from anomalies [4, 5]. This ideal is seldom met in practical scenarios, where the inclusion of even a few anomalous samples can severely skew the learning process, leading to either too restrictive or overly permissive models. Models trained under these conditions are prone to overfitting, resulting in many false positives during operational deployment. Conversely, including anomalies in the training set can escalate the incidence of false negatives, severely undermining the system’s reliability.
To illustrate the severity of this issue, Fig. 1 demonstrates a decline in AUCROC scores as noise increases in the training data. Traditional detection methods show a marked deterioration in performance as data imperfections escalate, highlighting their vulnerability. In contrast, our proposed Iterative Refinement Process (IRP) maintains a higher level of performance, showcasing its robustness and adaptability across a broad range of challenging conditions. We deliberately limit our evaluation to scenarios where the noise does not render the data overly corrupt, ensuring meaningful learning and generalization. This decision reflects our commitment to developing a method that optimally balances performance with practical applicability in realistically noisy environments.
@@
Addressing these complexities requires a robust detection system capable of navigating the intricacies of noisy and incomplete data sets. Current advancements in deep learning, such as autoencoders [6], generative adversarial networks (GANs) [7], and discrete feature space representations [8], offer promising ways for enhancing anomaly detection. However, these technologies still struggle with determining the optimal anomaly score function, ensuring sufficient robustness against noise and outliers, and generalizing effectively to novel, unseen data sets.
In response to these ongoing challenges, this paper introduces a methodology named Iterative Refinement Process (IRP) specifically designed for robust anomaly detection. This process combines advanced machine learning techniques with a cyclic data refinement strategy to iteratively enhance the training dataset’s quality, thereby significantly improving the model’s performance and robustness. IRP systematically identifies and removes the most misleading data points based on a dynamically adjusted scoring mechanism, thereby refining the model’s ability to generalize from normal operational data while minimizing the influence of outliers.
The main contributions of this paper are as follows:
  • We propose the Iterative Refinement Process (IRP), a robust training methodology that employs a cyclic refinement strategy to enhance data quality and model accuracy in anomaly detection systematically;
  • The dynamic threshold adjustment mechanism adapts the exclusion threshold based on a robust statistical measure, ensuring precise and adaptive outlier removal. This feature enhances the system’s ability to accurately identify and exclude data points that do not conform to the expected pattern, improving overall detection performance.
  • We present an experimental validation of our proposed methodology across two challenging public datasets, namely KSDD2 and MVTec-AD, demonstrating significant improvements over traditional approaches and establishing a new standard for robustness in the field.
@@ 2 Related Work
Detecting and identifying defects is crucial in manufacturing to ensure processes function correctly. Anomaly detection involves identifying issues such as scratches, blemishes, blockages, discoloration, holes, and, in general, any surface or structural irregularity on a manufactured item [9]. Recent surveys review AD’s state of the art from both technological and methodological perspectives [10, 11]. Traditionally, AD can be formulated as a supervised or unsupervised learning problem. Supervised methods use both anomaly and normal samples during training, while unsupervised methods only use normal samples at training time. While supervised methods achieve better results, the high cost of annotated datasets and advances in generative AI have shifted focus towards unsupervised approaches. Many unsupervised methods rely on image reconstruction using encoder-decoder networks to spot anomalies by comparing original and reconstructed images [12, 13, 14]. These networks struggle to reconstruct anomalous regions accurately, having never encountered them during training, which can result in high false positive rates with variable training data. To address this, [15] employs a reconstruction network to restore normal appearances in anomalous images, while another approach introduces textual prompts describing defects as in [16, 17]. Comparing features instead of whole images also helps to reduce false positives [6, 18].
Robust anomaly detection (RAD) addresses the problem of mitigating the effect of bad annotations in the training set. It has been widely studied in several fields, using methods ranging from robust statistical techniques [19] to deep learning approaches like autoencoders [5, 20] and recurrent neural networks [21]. These methods all aim to reduce sensitivity to noise in the labeling process. Alternatively, [22] focused on removing noisy labels while using high-performing anomaly detectors. Here, a two-layer online learning framework filters suspicious data and detects anomalies from the remaining data. We argue that training two different models (one for predicting data quality and one for anomaly detection) adds unnecessary complexity. More recently, researchers have explored self-supervised learning (SSL). Most of the works rely on adaptations of deep CNN for enhancing learning capabilities from unlabelled data  [23, 24, 25], or using domain-adaptation techniques [26]. In our previous work [27], we showed that SSL could also be effective in filtering badly labeled data in the training set. In this paper, we extend the previous work by using an iterative approach that is based on statistical observations of anomaly score estimation on the training set.
@@ 3 Iterative Refinement Process
The Iterative Refinement Process (IRP) introduces a novel approach to significantly enhance the robustness and accuracy of anomaly detection in an industrial environment by applying advanced probabilistic models to iteratively refine the training dataset. Despite being agnostic with respect to the anomaly detection model, for our preliminary evaluation of the impact of our approach, we used the DifferNet model [18], which provided foundational insights into the potential refinements necessary for handling complex anomaly detection in industrial settings. The DifferNet model employs a normalizing flow-based architecture to provide precise density estimations of image features extracted via a convolutional neural network. DifferNet establishes a bijective relationship between the feature space and the latent space, enabling each vector to be uniquely mapped to a likelihood score. This scoring is derived from the model’s ability to discern common patterns from uncommon ones, making it especially effective for detecting subtle anomalies often found in defect detection scenarios. Initially, each data point, $x_{i}$, is evaluated using a probabilistic model tailored to anomaly detection, calculating an anomaly likelihood as follows:
@@
where $p(x_{i};\theta)$ is the probability density function of data points with the model parametrization of $\theta$. The framework engages in a dynamic data refinement cycle, where each point is assessed against an adaptive threshold calculated as $T$ times the median of the anomaly scores, $A(x_{i})$. If any data point’s anomaly score exceeds this threshold, it is removed to refine the dataset. The refined dataset, denoted $D_{\text{new}}$, excludes the most anomalous data point, $x_{\text{max}}$, where $x_{\text{max}}$ is the point with the maximum anomaly score exceeding the threshold. This process of refinement and retraining on $D_{\text{new}}$ continues until the dataset achieves optimal stability and performance. The steps involved in this refinement process are elaborated in Algorithm 1, and the entire sequence from data input through anomaly detection, refinement, and final evaluation is illustrated in Fig. 2.
Algorithm 1 Iterative Refinement and Retraining
1:Input: Dataset of images $D=\{x_{1},x_{2},\ldots,x_{n}\}$
2:Pre-train model on $D$ for $N$ epochs
3:Compute anomaly scores $A=\{a_{1},a_{2},\ldots,a_{n}\}$ where $a_{i}=-\log p(x_{i};\theta)$
4:Compute median of scores $M=\text{median}(A)$triangleright Median score used for thresholding
5:if $\max(A)>T\times M$ then
6:     $x_{\text{max}}\leftarrow x_{i}$ where $a_{i}=\max(A)$
7:     Delete $x_{\text{max}}$ from $D$
8:     $D_{\text{new}}\leftarrow D\setminus\{x_{\text{max}}\}$
9:else
10:     $D_{\text{new}}\leftarrow D$
11:end if
12:Retrain model on $D_{\text{new}}$
13:repeat
14:     Repeat steps 3 to 9 until the specified number of epochs is reached
15:until Maximum epochs reached
@@ 3.1 Probabilistic Anomaly Scoring Mechanism and Dynamic Refinement
Anomaly scores are computed directly from the latent space using the likelihood of the transformed features. Lower likelihood values indicate higher probabilities of an anomaly due to the deviation of features from the model of ’normal’ data. These scores are calculated by averaging the negative log-likelihoods of multiple transformed instances of an image $x$, which are obtained using a set of predefined transformations $T$:
@@
Our significant contribution is manifest in the dynamic adjustment of the assessment threshold $\theta$, which is iteratively recalibrated in response to evolving data characteristics and directly influenced by the statistical nature of the anomaly scores:
$\text{Threshold}=T\times M$ (5)
Data points whose anomaly scores exceed this threshold are considered outliers and are subsequently removed from the dataset:
@@ 4.1 Datasets
Kolektor SDD2 (KSDD2): The KSDD2 dataset [2] consists of RGB images showcasing defective production items, meticulously sourced and annotated by Kolektor Group d.o.o. The defects vary widely in size, shape, and color, ranging from minor scratches and spots to significant surface faults. For uniform evaluation, all images are center-cropped and resized to dimensions of 448 x 448 pixels. The training set comprises 2085 normal samples and 246 positive samples, while the testing set includes 894 negative samples and 110 positive samples.
MVTec-AD Dataset: The MVTec-AD dataset [28] is utilized to demonstrate the effectiveness of our proposed method across a variety of real-world industrial products. The MVTec-AD dataset contains 15 different types of industrial products, encompassing over 70 different types of defects, each labelled with both defect types and segmentation masks. To address the original lack of defective images in the training set, we resampled half of the defect images to include in our training data. The revised training set contains 889 normal and 1345 defective images, while the testing set comprises 1210 normal and 724 defective images.
@@ 4.2 Implementation Details
Experiments were conducted using PyTorch on NVIDIA GPUs, specifically utilizing CUDA. Our model processes images resized to $448\times 448$ pixels, with standardized preprocessing that includes optional rotations and normalization using means of [0.485, 0.456, 0.406] and standard deviations of [0.229, 0.224, 0.225]. The network architecture features three scales of input resolution, 8 coupling blocks within the normalizing flow model, and fully connected layers in the scale-translation networks each containing 2048 neurons, with no dropout our proposed method’s effectiveness across variousapplied. Training parameters are set with a learning rate of $2\times 10^{-4}$, and the model undergoes 400 epochs with batch size 96. Data augmentation involves multiple transformations per sample, with 4 in training and 64 in testing to ensure robustness. Outputs are managed with verbose logging for detailed monitoring, optional gradient map visualization, and periodic model saving to capture progress and facilitate ongoing optimization.
4.3 Varying Noise Levels on KSDD2

### v2

# Self-Supervised Iterative Refinement for Anomaly Detection in Industrial Quality Control

@@ Abstract
This study introduces the Self-Supervised Iterative Refinement Process (IRP), a robust anomaly detection methodology tailored for high-stakes industrial quality control. The IRP leverages self-supervised learning to improve defect detection accuracy by employing a cyclic data refinement strategy that iteratively removes misleading data points, thereby improving model performance and robustness. We validate the effectiveness of the IRP using two benchmark datasets, Kolektor SDD2 (KSDD2) and MVTec-AD, covering a wide range of industrial products and defect types. Our experimental results demonstrate that the IRP consistently outperforms traditional anomaly detection models, particularly in environments with high noise levels. This study highlights the potential of IRP to significantly enhance anomaly detection processes in industrial settings, effectively managing the challenges of sparse and noisy data.
@@ 1 INTRODUCTION
Anomaly detection (AD) plays an indispensable role in quality control in a wide range of manufacturing industries, ensuring the integrity of materials such as marble [Vrochidou et al., 2022], steel [Božič et al., 2021b], and leather [Jawahar et al., 2023]. Accurately detecting items that are non-compliant with product specifications is crucial for maintaining product standards and consumer satisfaction. However, this task is fraught with challenges, primarily due to the diverse and complex textures of the materials, the rarity of defects, and the significant scarcity of accurately labeled data necessary for effective supervised learning. Anomalies typically occupy only a tiny portion of an image, making their detection difficult for human inspectors and automated systems. This challenge is compounded by the labor intensive and error-prone nature of manually labeled training data, a process that becomes increasingly unsustainable in high-throughput manufacturing settings.
The evolution of anomaly detection in recent years has increasingly leaned towards unsupervised and self-supervised learning paradigms, prized for their ability to operate without extensively labeled datasets. These approaches, while innovative, presuppose the existence of pristine training data, free from anomalies [Ono et al., 2020, Beggel et al., 2020]. This ideal is seldom met in practical scenarios, where the inclusion of even a few anomalous samples can severely skew the learning process, leading to either too restrictive or overly permissive models. Models trained under these conditions are prone to overfitting, resulting in many false positives during operational deployment. In contrast, including anomalies in the training set can escalate the incidence of false negatives, severely undermining the system’s reliability.
@@
To illustrate the severity of this issue, Fig. 1 demonstrates a decrease in AUCROC scores as noise increases in the training data. Traditional detection methods show a marked deterioration in performance as data imperfections increase, highlighting their vulnerability. In contrast, our proposed Self-Supervised Iterative Refinement Process (IRP) maintains a higher level of performance, showcasing its robustness and adaptability across a broad range of challenging conditions. We deliberately limit our evaluation to scenarios where the noise does not render the data overly corrupt, ensuring meaningful learning and generalization. This decision reflects our commitment to developing a method that optimally balances performance with practical applicability in realistically noisy environments.
Addressing these complexities requires a robust detection system capable of navigating the intricacies of noisy and incomplete data sets. Current advances in deep learning, such as autoencoders [Zavrtanik et al., 2022], generative adversarial networks (GAN) [Luo et al., 2022], and discrete feature space representations [Hu et al., 2022], offer promising ways to improve anomaly detection. However, these technologies still struggle with determining the optimal anomaly score function, ensuring sufficient robustness against noise and outliers, and generalizing effectively to novel, unseen data sets.
In response to these ongoing challenges, this paper introduces a methodology named Self-Supervised Iterative Refinement Process (IRP) specifically designed for robust anomaly detection. This process combines advanced machine learning techniques with a cyclic data refinement strategy to iteratively enhance the training data set quality, thereby significantly improving the model’s performance and robustness. IRP systematically identifies and removes the most misleading data points based on a dynamically adjusted scoring mechanism, thereby refining the model’s ability to generalize from normal operational data while minimizing the influence of outliers.
The main contributions of this paper are as follows.
  • We propose the Self-Supervised Iterative Refinement Process (IRP), a robust training methodology that employs a cyclic refinement strategy to enhance data quality and model accuracy in anomaly detection systematically;
  • The dynamic threshold adjustment mechanism adapts the exclusion threshold based on a robust statistical measure, ensuring precise and adaptive outlier removal. This feature improves the system’s ability to accurately identify and exclude data points that do not conform to the expected pattern, improving overall detection performance.
  • We present an experimental validation of our proposed methodology in two challenging public datasets, namely KSDD2 and MVTec-AD, which demonstrate significant improvements over traditional approaches and establish a new standard for robustness in the field.
@@ 2 RELATED WORK
Detecting and identifying defects is crucial in manufacturing to ensure processes function correctly. Anomaly detection involves identifying issues such as scratches, blemishes, blockages, discoloration, holes, and, in general, any surface or structural irregularity on a manufactured item [De Vitis et al., 2020]. Recent surveys review AD’s state of the art from both technological and methodological perspectives [Chen et al., 2021, Bhatt et al., 2021]. Traditionally, AD can be formulated as a supervised or unsupervised learning problem. Supervised methods use both anomaly and normal samples during training, while unsupervised methods only use normal samples at training time. While supervised methods achieve better results, the high cost of annotated datasets and advances in generative AI have shifted focus towards unsupervised approaches. Many unsupervised methods rely on image reconstruction using encoder-decoder networks to spot anomalies by comparing original and reconstructed images [Akcay et al., 2019, Defard et al., 2021, Roth et al., 2022]. These networks struggle to reconstruct anomalous regions accurately, having never encountered them during training, which can result in high false positive rates with variable training data. To address this, [Zavrtanik et al., 2021] employs a reconstruction network to restore normal appearances in anomalous images, while another approach introduces textual prompts describing defects as in [Capogrosso et al., 2024, Girella et al., 2024]. Comparing features instead of whole images also helps to reduce false positives [Zavrtanik et al., 2022, Rudolph et al., 2021].
Robust anomaly detection (RAD) addresses the problem of mitigating the effect of bad annotations in the training set. It has been widely studied in several fields, using methods ranging from robust statistical techniques [Rousseeuw and Hubert, 2018] to deep learning approaches like autoencoders [Beggel et al., 2020, Zhou and Paffenroth, 2017] and recurrent neural networks [Su et al., 2019]. These methods all aim to reduce sensitivity to noise in the labeling process. Alternatively, [Zhao et al., 2019] focused on removing noisy labels while using high-performing anomaly detectors. Here, a two-layer online learning framework filters suspicious data and detects anomalies from the remaining data. We argue that training two different models (one for predicting data quality and one for anomaly detection) adds unnecessary complexity. More recently, researchers have explored self-supervised learning (SSL) as a promising avenue for anomaly detection. SSL methods often rely on adaptations of deep CNN for enhancing learning capabilities from unlabelled data  [Božič et al., 2021a, Zhang et al., 2022, Tian and Jia, 2022], or using domain-adaptation techniques [Zhang et al., 2021]. In our previous work [Aqeel et al., 2024], we demonstrated that SSL could be effective in filtering badly labeled data in the training set. This paper extends our previous work by using an iterative refinement approach, which is based on statistical observations of anomaly score estimation on the training set, to further improve robustness and accuracy in anomaly detection.
@@ 3 ITERATIVE REFINEMENT PROCESS
The Iterative Refinement Process (IRP) introduces a novel approach to significantly enhance the robustness and accuracy of anomaly detection in an industrial environment by applying advanced probabilistic models to iteratively refine the training dataset. Despite being agnostic with respect to the anomaly detection model, for our preliminary evaluation of the impact of our approach, we used the DifferNet model [Rudolph et al., 2021], which provided foundational insights into the potential refinements necessary for handling complex anomaly detection in industrial settings. The DifferNet model employs a normalizing flow-based architecture to provide precise density estimations of image features extracted via a convolutional neural network. DifferNet establishes a bijective relationship between the feature space and the latent space, enabling each vector to be uniquely mapped to a likelihood score. This scoring is derived from the model’s ability to discern common patterns from uncommon ones, making it especially effective for detecting subtle anomalies often found in defect detection scenarios. Initially, each data point $x_{i}$, which refers to a transformed image, is evaluated using a probabilistic model tailored to anomaly detection. The model calculates an anomaly likelihood as follows:
@@
where $p(x_{i};\theta)$ is the probability density function of the transformed image $x_{i}$, with $\theta$ representing the parameters of the model. The framework engages in a dynamic data refinement cycle, where each point is assessed against an adaptive threshold calculated as $\lambda$ times the median of the anomaly scores, $A(x_{i})$. If any data point’s anomaly score exceeds this threshold, it is removed to refine the dataset. The refined dataset, denoted $D_{\text{new}}$, excludes the most anomalous data point, $x_{\text{max}}$, where $x_{\text{max}}$ is the point with the maximum anomaly score exceeding the threshold. This process of refinement and retraining on $D_{\text{new}}$ continues until the dataset achieves optimal stability and performance. The steps involved in this refinement process are elaborated in Algorithm 1, and the entire sequence from data input through anomaly detection, refinement, and final evaluation is illustrated in Fig. 2.
@@ 3.1 Probabilistic Anomaly Scoring Mechanism and Dynamic Refinement
The anomaly scores are derived directly from the latent space by evaluating the likelihood of the transformed features. A lower likelihood indicates a higher probability of an anomaly, as it reflects the deviation of the features from the model of ’normal’ data. These scores are computed by averaging the negative log-likelihoods of multiple transformed versions of an image $x$, which are generated using a set of predefined transformations $T$. The transformations $T$ represent a collection of operations that perturb or modify the original data in controlled ways to simulate variations that may highlight anomalies more clearly. These transformations may include geometric modifications, such as rotations, translations, and scalings, or they may involve noise injections or other perturbations that simulate real-world variations in the data.
The purpose of applying these transformations is twofold. First, it helps with data augmentation by generating multiple alternative versions of the input data, thereby making the anomaly detection process more robust to specific features of the original data. Second, it stabilizes the anomaly scoring by reducing the impact of random noise introduced by any single transformation. By averaging the scores over several transformed instances of the data, the final anomaly score becomes more reliable, reflecting the underlying anomaly rather than spurious noise.
For each transformation $T_{i}\in T$, the anomaly score is calculated by applying $T_{i}$ to the input data $x$, passing it through the feature extractor $f_{ex}$, and then mapping the resulting features to the latent space using the normalizing flow model $f_{NF}$. The final score is the expected value of the negative log-likelihoods of these transformed instances, as shown in Equation 4. This approach ensures that the anomaly score reflects the data’s likelihood in the transformed space, improving the robustness of the anomaly detection process by incorporating multiple views of the data.
@@
Our significant contribution is manifest in the dynamic adjustment of the assessment threshold $\theta$, which is iteratively recalibrated in response to evolving data characteristics and directly influenced by the statistical nature of the anomaly scores. We chose the median as a central measure for thresholding due to its robustness against outliers and skewed data distributions. The median offers a more stable and representative central tendency, particularly in datasets where the majority of values are concentrated around a certain range with a few extreme outliers. This characteristic makes the median especially suitable for determining thresholds in anomaly detection, where the presence of anomalies can distort the distribution of scores. The threshold defined as:
$\text{Threshold}=\lambda\times M$ (5)
where $\lambda$ is the scaling factor and $M$ is the median of the anomaly scores. This ensures that data points with anomaly scores exceeding the threshold are considered outliers and are removed from the dataset.
@@
Dataset of images $D=\{x_{1},x_{2},\ldots,x_{n}\}$ Pre-train model on $D$ for $N$ epochs
Compute anomaly scores $A=\{a_{1},a_{2},\ldots,a_{n}\}$ where $a_{i}=-\log p(x_{i};\theta)$
Compute median of scores $M=\text{median}(A)$
if $\max(A)>\lambda\times M$ then
       $x_{\text{max}}\leftarrow x_{i}$ where $a_{i}=\max(A)$
       Delete $x_{\text{max}}$ from $D$
       $D_{\text{new}}\leftarrow D\setminus\{x_{\text{max}}\}$
else
       $D_{\text{new}}\leftarrow D$
Retrain model on $D_{\text{new}}$
repeat
       Repeat steps 3 to 9
until Convergence
Algorithm 1 Iterative Refinement and Retraining
@@ 4.1 Datasets
Kolektor SDD2 (KSDD2): The KSDD2 dataset [Božič et al., 2021b] consists of RGB images showcasing defective production items, meticulously sourced and annotated by Kolektor Group d.o.o. The defects vary widely in size, shape, and color, ranging from minor scratches and spots to significant surface faults. For uniform evaluation, all images are center-cropped and resized to dimensions of 448$\times$448 pixels. The training set comprises 2085 normal and 246 positive samples, while the testing set includes 894 negative and 110 positive samples.
MVTec-AD Dataset: The MVTec-AD dataset [Bergmann et al., 2019] is utilized to demonstrate the effectiveness of our proposed method across a variety of real-world industrial products. The MVTec-AD dataset contains 15 different types of industrial products, encompassing over 70 different types of defects, each labelled with both defect types and segmentation masks. To address the original lack of defective images in the training set, we resampled half of the defect images to include in our training data. The revised training set contains 889 normal and 1345 defective images, while the testing set comprises 1210 normal and 724 defective images.
[KSDD2]Refer to caption [MVTec-AD]Refer to caption
@@ 4.2 Implementation Details
Experiments were conducted using PyTorch framework. We resized images to $448\times 448$ pixels and applied standardized preprocessing techniques, including optional rotations and normalization. The network architecture features three scales of input resolution, 8 coupling blocks within the normalizing flow model, and fully connected layers in the scale-translation networks each containing 2048 neurons, with no dropout our proposed method’s effectiveness across variousapplied. Training parameters are set with a learning rate of $2\times 10^{-4}$, and the model undergoes 400 epochs with batch size 96.
We compared our proposed methodology with existing models including DifferNet [Rudolph et al., 2021] and One Shot Removal (OSR) [Aqeel et al., 2024]. DifferNet leverages a multi-scale feature extraction process to assign meaningful likelihoods to images, thereby facilitating defect localization. DifferNet represents also the model for feature extraction used in our self supervised procedure. OSR enhances the robustness of surface defect detection models through a novel training pipeline, which involves initial training, anomalous sample removal, and model fine-tuning.
The performance of our model was tested under varying noise levels, from 0% to 50%. These noise levels simulate the inclusion of bad samples typically encountered in unsupervised learning scenarios. We measured the model’s performance using the AUROC score across these noise levels, providing a reliable indicator of its ability to distinguish between normal and anomalous data. Each experiment was repeated three times to ensure consistency and to assess the impact of noise on detection accuracy.
4.3 Performance on KSDD2 Across Varying Noise Levels